If the price of a bottle of orange juice is $0.75, the total consumer surplus received by these consumers is S (Enter your response rounded to hwo decimal places) Suppose the price of a bottle of orange juice rises to $150 Once the price of a bottle of orange juice rises to $1 50, the total consumer surplus received by these consumers is : $(Enter your response rounded to hwo decimal places) If the price of a bottle of orange juice is 50.75,then consumer surplus is an amount equal to the area OA above the demand curve for each individual who buys and consumes orange juice OB. under the demand curve for each individual who buys and consumes orange juice OC. under the demand curve and above the 50 75 price for each individual whe buys and consumes orange juice OD. under the demand curve and above the $1 50 price for each individual whe buys and consumes orange juice If the price of a botte of orange juice is $1.50, then consumer surplus is an armount equal to the area OA above the demand curve for each individual who buys and consumes orange juice OB. under the demand curve and above the $1.50 price for each individual whe buys and consumes orange luice OC under the demand curve for each individual who buys and consumes orange juice OD. under he demand curve and above the 50 75 price for each individual whe buys and consumes onange jice
n the market for orange juice. Highest Price Willing to Pay $4 Consumer Jill Jose 2 Josh Jordan $(Enter your response rounded to two decimal places.) mer surplus received by these consumers is Cinun nnaanan sasundad tn hain danimal nlanee
JUIUar If the price of a bottle of orange juice is $0.75, the total consumer surplus received by these consumers is S (Enter your respon Suppose the price of a bottle of orange juice rises to $1.50. Once the price of a bottle of orange juice rises to $1.50, the total consumer surplus received by these consum (Enter yoe umers is $ If the price of a bottle of orange juice is $0.75, then consumer surplus is an amount equal to the area OA. above the demand curve for each individual who buys and consumes orange juice O B. under the demand curve for each individual who buys and consumes orange juice. OC. under the demand curve and above the $0.75 price for each individual who buys and consumes orange juice. O D. under the demand curve and above the $1.50 price for each individual who buys and consumes orange juice. If the price of a bottle of orange juice is $1.50, then consumer surplus is an amount equal to the area OA. above the demand curve for each individual who buys and consumes orange juice O B. under the demand curve and above the $1.50 price for each individual who buys and consumes orange juice. OC. under the demand curve for each individual who buys and consumes orange juice O D. under the demand curve and above the $0.75 price for each individual who buys and consumes orange juice. Click to seant u ne ls
Answer #1


Ans A)

Total Consumer Surplus= Sum Of Consumer surplus for each consumer in the market

We have 4 consumers

Individual Consumer Surplus= Willingness to pay of Individual-Market Price

Total Consumer Surplus=Sum of Willingness to Pay for all Consumers-(Number of consumers)*Market Price

=10-4(Market Price)

Ans B)

If Price of a Bottle of Orange Juice is $0.75 then Consumer Surplus=(4+3+2+1)-(4*0.75)=$7

Ans C)

If Price of a Bottle of Orange Juice is $1.50 then Consumer Surplus=(4+3+2+1)-(4*1.5)=$4


Ans D)
If the price of bottle of Orange Juice is $0.75 then Total consumer surplus is the area under demand curve and above Market Price level $0.75

Hence option C is correct graphical representation of Consumer Surplus

Ans E)
If the price of bottle of Orange Juice is $1.50 then Total consumer surplus is the area under demand curve and above Market Price level $1.50

Hence option B is correct graphical representation of Consumer Surplus